# Powering In-Database Dynamic Model Slicing for Structured Data Analytics

### v1

@@ 4. SQL-Aware Dynamic Model Slicing
In contrast to conventional machine learning paradigms, SQL-aware predictive modeling makes use of constraints specified in SQL queries to provide more accurate predictions with respect to the data of interest. For example, the query outlined in Figure 2 is interested in data constrained based on age, location, and gender. Such selection constraints present optimization opportunities for prediction accuracy and efficiency. To this end, we propose SQL-aware dynamic model slicing technique, LEADS, to leverage the propositional formula $\varphi$ from SQL queries as meta-information to customize the base model, as a means to improve the effectiveness and efficiency of the prediction of a specified datasubset.
In this section, we first introduce the SQL query encoder to translate $\varphi$ into a vectorized format. Next, we present two key components of LEADS for scaling up the modeling capacity of the base model via the Mixture of Experts (MoE) technique and dynamic model slicing via a SQL-aware gating network. For optimization, we further design two regularization terms to strike a balance between effectiveness and efficiency.
@@
(b) Process of Encoding SQL query.
@@
4.1. SQL Query Encoder
In SQL-aware predictive modeling, the WHERE clause of SQL queries filters tuples according to a predefined propositional formula $\varphi$. This formula comprises one or more predicates, each setting a logical condition on a particular attribute. For instance, ”gender = ‘M’ ” mandates that the gender attribute of the filtered tuples must be ‘M’. These predicates are interconnected via logical operators such as ”AND” or ”OR” to form a complete propositional formula. The exponential number of possible predicate combinations in SQL queries renders the direct development of models for every conceivable subdataset, as in traditional machine learning approaches, both complex and intractable.
For the SQL query encoder, we focus on individual queries, referred to as primitive SQL query. Considering a table T with $M$ attributes, the $j$-th attribute denoted as $A_{j}$, each attribute is linked to either a numerical or categorical feature in predictive modeling. Particularly, each numerical feature needs to be converted into a corresponding categorical feature through discretization, which will be detailed subsequently. In a primitive SQL query, each attribute $A_{j}$ may be associated with zero or one predicate, with predicates across attributes conjoined using the logical operator $\land$ (AND), as depicted in Figure 2(a). Technically, a predicate for attribute $A_{j}$ in primitive SQL query can be expressed as $P_{j}:A_{j}=a_{j}$, where $a_{j}\in\mathcal{D}_{j}\cup\{\Delta_{j}\}$, $\mathcal{D}_{j}$ represents the domain of possible values for $A_{j}$, and $\Delta_{j}$ denotes a default value assigned to $A_{j}$ when it is not specified in the query. Figure 2(a) illustrates a valid primitive SQL query example, contrasting with two non-examples. Thus, the propositional formula $\varphi$ can be represented as:
$\varphi=P_{1}\land P_{2}\land\cdots\land P_{M}.$
The objective of the SQL query encoder is to generate a categorical feature vector $\mathbf{q}$ for each primitive SQL query based on the meta-information $\varphi$, achieved by concatenating the attribute values of the predicates. Formally, the feature vector of the SQL query encoding can be obtained by:
$\mathbf{q}=[q_{1},q_{2},\cdots,q_{M}]$
where $q_{j}$ is the categorical attribute value for predicate $P_{j}$. Figure 2(b) demonstrates the transformation of a primitive SQL query into a feature vector. Notably, the numerical attribute ”age” here is first discretized before being encoded alongside categorical attributes ”city” and ”gender”, and columns lacking predicates are filled by the default value $\Delta_{i}$.
@@ 4.1. SQL Query Encoder
Discretization. Discretization is essential for encoding numerical attributes, e.g., weight or salary. The infinite possible values of numerical attributes make direct encoding infeasible, hence requiring discretization. This process first partitions the domain $\mathcal{D}$ of each numerical attribute into a fixed number of bins, akin to approximating a k-nearest neighbors classifier in predictive modeling. The goal of discretization is to preserve the key information in the embedding space for maintaining predictive modeling capacity.
To this end, we employ a supervised discretization approach that accounts for the correlation between numerical attributes and the target attribute. This aims to maximize information value (IV), which measures the reduction of uncertainty within each bin relative to the prediction target. Higher IV values indicate a significant decrease in uncertainty, thereby preserving the predictive capacity. In particular, we introduce the open-source OptBinning (Navas-Palencia, 2020) implementation for discretization, which optimizes IV effectively while supporting constraints like the maximum bin count per attribute.
4.2. SQL-Aware Dynamic Model Slicing
The categorical feature vector q, obtained from the SQL query encoder, captures key information that facilitates dynamic customization of a predictive model to an optimal configuration for the targeted subdataset. The customization should significantly enhance predictive performance in SQL-aware predictive modeling. As illustrated in Figure 3, LEADS first scales up the modeling capacity via MoE by replicating the base model to construct a general model, and subsequently, LEADS integrates a SQL-aware gating network based on the SQL encoding vector q to selectively activate experts in the general model to derive a sliced model for higher efficiency and effectiveness.
In this subsection, we will first introduce the preprocessing module that prepares the embedding vector q for the predictive modeling, elaborate on the two key modules, i.e., the general model and the SQL-aware gating network, and finally, explain our SQL-aware dynamic model slicing technique in detail.
4.2.1. Preprocessing Module
There are two sets of input constructed for the SQL-aware prediction modeling given an input tuple. The first set of input is constructed for the gating network and can be uniformly represented as a categorical feature vector $\mathbf{q}$. $\mathbf{q}=[q_{1},q_{2},\ldots,q_{M}]$ comprises $M$ feature values from respective attribute fields, where numerical attributes need to be converted into categorical attributes via discretization, as discussed in the previous subsection. The second set is the attribute values of the input tuple $\mathbf{x}=[x_{1},x_{2},\ldots,x_{M}]$, and each attribute value $x_{i}$ can be either categorical or numerical.
For both $\mathbf{q}$ and $\mathbf{x}$, each field of attribute value $v_{i}$ ($q_{i}$/$x_{i}$) needs to be transformed into a corresponding embedding vector $\mathbf{e}_{i}$ to participate the subsequent predictive modeling. Specifically, each categorical attribute is transformed via embedding lookup, i.e., $\mathbf{e}_{i}=\mathbf{E}_{i}[q_{i}],\mathbf{e}_{i}\in\mathbb{R}^{n_{e}}$, where $n_{e}$ is the feature embedding size, and $\mathbf{E}_{i}$ is the embedding matrix of this categorical attribute. Note that different embedding vectors of $\mathbf{E}_{i}$ correspond to their respective values of this attribute. As for each numerical attribute $x_{j}$ of $\mathbf{x}$, the corresponding embedding vector is obtained by linearly scaling up a learnable embedding vector $\mathbf{\hat{e}}_{j}$ for this numerical attribute, namely $\mathbf{e}_{j}=x_{j}\cdot\mathbf{\hat{e}}_{j}$. In this way, we obtain fixed-size inputs, i.e., embedding vectors $\mathbf{\hat{q}}=[\mathbf{q}_{1},\mathbf{q}_{2},\ldots,\mathbf{q}_{M}]$ and $\mathbf{\hat{x}}=[\mathbf{x}_{1},\mathbf{x}_{2},\ldots,\mathbf{x}_{M}]$.
4.2.2. General Model and SQL-aware Gating Network
@@
The general model comprises a set of $K$ replicated base models, denoted as $\mathbf{\mathcal{F}}=[\mathcal{F}_{1},\mathcal{F}_{2},\ldots,\mathcal{F}_{K}]$, which are referred to as ”expert models”. These experts share the same model architecture but learn distinct model parameters during training, which take the same input $\mathbf{\hat{x}}$ and produce different outputs that need to be aggregated for final predictions. The output of the $i$-th expert for a given input $\mathbf{x}$ is denoted as $\mathcal{F}_{i}(\mathbf{\hat{x}})$.
@@
As for the SQL-aware gating network $\mathcal{G}$, it takes the SQL query embedding vectors $\mathbf{\hat{q}}$ as input to produce a $K$-dimensional vector, termed the gating weight $\mathbf{w}$, with $\mathbf{w}\in\mathbb{R}^{K}$. Specifically, a two-layer multilayer perceptron (MLP) is employed as the gating network following the practice (Riquelme et al., 2021; Fedus et al., 2022; OpenAI, 2023). We concatenate all embeddings in $\mathbf{\hat{q}}$ as the input of the gating network $\widetilde{\textbf{q}}=\mathbf{q}_{1}\oplus\mathbf{q}_{2}\ldots\oplus\mathbf{q}_{M}$, where $\widetilde{\textbf{q}}\in\mathbb{R}^{M{\cdot}{n_{e}}}$, then feed $\widetilde{\mathbf{q}}$ to $\mathcal{G}$, and obtain the gating weight $\mathbf{w}$ by:
@@
$\mathbf{w}$ $=G(\mathbf{s})=\mathbf{W_{2}}\mathbf{z}+\mathbf{b_{2}}$
@@
Given the gating weight $\mathbf{w}$, the $\alpha$-entmax function (Peters et al., 2019; Correia et al., 2019) is further applied to recalibrate $\mathbf{w}$ to a probability distribution. As introduced in Section 2, the hyper-parameter $\alpha$ in $\alpha$-entmax controls the level of sparsity, and a larger value of $\alpha$ sets more gating weights to zero and thus deactivates more experts for higher efficiency. The output of $\alpha$-entmax $\widetilde{\mathbf{w}}$ is thus:
@@
which is used to aggregate expert outputs. The final output of the general model is a weighted average of expert outputs:
@@
where $\mathbf{\hat{y}}$ is the prediction given the input $\mathbf{x}$ and the corresponding query $\mathbf{q}$ of the SQL-aware predictive modeling.
4.2.3. Dynamic Model Slicing via Gating Network
For a given SQL query, all the retrieved data tuples share the same recalibrated gating weight $\widetilde{\mathbf{w}}$. Further, $\widetilde{w}_{i}=0$ in Equation 5 indicates that the corresponding $i$-th expert is not required in the current predictive modeling, and thus, only a small fraction of experts $\mathcal{F}_{i}$ need to be activated for prediction for much higher computational efficiency.
Denoting the set of indices of activated experts as $\{I_{1},I_{2},\cdots,I_{n_{o}}\}$, where $n_{o}$ is the current number of activated experts and $\widetilde{w}_{I_{j}}\neq 0,\forall j\in\{1,2,\dots,n_{o}\}$, and given the corresponding SQL query encoding $\mathbf{q}$, we index the activated experts to form a sliced model, i.e., $\mathbf{\mathcal{F}}_{\mathbf{q}}=[\mathcal{F}_{I_{1}},\mathcal{F}_{I_{2}},\cdots,\mathcal{F}_{I_{n_{o}}}]$. Therefore, the final output of the sliced model is as follows:
@@ 4.2.3. Dynamic Model Slicing via Gating Network
where the number of activated experts $n_{o}$ directly affects the effectiveness and efficiency of the sliced model. A large $n_{o}$ indicates larger model capacity while incurring higher computational overhead, and vice versa. In LEADS, $n_{o}$ is determined by the gating network based on the SQL query encoding q and the hyper-parameter $\alpha$ of the sparse softmax function. Notably, instead of predefining a fixed value, $\alpha$ in $\alpha$-entmax is learnable and optimized based on the input tuples and corresponding queries during training. Subsequently, during inference, LEADS can dynamically adapt $n_{o}$ based on the current SQL query, trading off between the effectiveness and efficiency of the predictive modeling.
@@ 4.3. Optimization
Our LEADS framework can be applied to different predictive tasks by configuring a proper objective function, such as mean squared error (MSE) for regression or cross-entropy for classification. For instance, in binary classification, the objective function employed is binary cross-entropy:
(7) ${\rm LogLoss}({\hat{\mathbf{y}},\mathbf{y}})=-\frac{1}{N}\sum_{i}^{N}\{y_{i}{\rm log}\sigma(\hat{y_{i}})+(1-y_{i}){\rm log}(1-\sigma(\hat{y_{i}}))\}$
where $\mathbf{\hat{y}}$ represents the prediction labels, $\mathbf{y}$ denotes the ground truth labels, $N$ is the number of tuples for prediction, and $\sigma(\cdot)$ is the sigmoid function.
To make the optimization more robust and effective, we introduce two regularization terms to the main loss function. The first term is the balance loss, $\mathcal{L}_{baln}$, which is introduced to address the issue of imbalanced expert utilization. This imbalance occurs when the gating network $\mathcal{G}$ tends to favor a small subset of experts, leading to a skewed training process where these preferred experts are overutilized while others are underutilized. Such a scenario undermines the capacity of MoE and can detrimentally affect the model performance.
Let $\mathbf{X}$ denote a mini-batch of training instances with $n_{b}$ tuples, and $\widetilde{\mathbf{W}}=[\widetilde{\mathbf{w}}_{1},\widetilde{\mathbf{w}}_{2},\cdots,\widetilde{\mathbf{w}}_{n_{b}}]$ is the recalibrated gating weights of $\mathbf{X}$, where $\widetilde{w}_{ij}$ is the $j$-th weight of $\widetilde{\mathbf{w}}_{i}$. $\mathcal{L}_{baln}$ is defined as:
@@
where $\mathbb{E}(\mathbf{\Phi})=\frac{1}{K}\sum_{j=1}^{K}\phi_{j}$. The balance loss term $L_{baln}$ encourages uniform distribution of weights across experts within the mini-batch to ensure more balanced importance among all experts.
The balance loss term is designed to encourage a more balanced selection of experts, which however, empirically results in activating a large number of experts despite the introduction of sparse softmax. To maintain sparsity within the model and counteract this tendency for more efficient computation, we further introduce a sparsity loss term, $\mathcal{L}_{sprs}$:
@@
which encourages the gating network to allocate higher weights to select a few experts while minimal or zero weights to others. Both loss terms are scaled by their respective regularization coefficient, $\lambda_{1}$ and $\lambda_{2}$, and then added to the main loss:
(10) $Loss={\rm LogLoss}(\hat{\mathbf{y}},{\mathbf{y}})+\lambda_{1}L_{baln}+\lambda_{2}L_{sprs}.$
With this objective function, LEADS can then be trained effectively with gradient-based optimizers, e.g., SGD or Adam (Kingma and Ba, 2014).

### v2

@@ 3. SQL-Aware Dynamic Model Slicing
In this section, we introduce our SQL-aware dynamic model slicing technique, LEADS, for supporting in-database predictive analytics within existing RDBMSs such as PostgreSQL. Unlike conventional machine learning approaches that make predictions based solely on features of individual instances, LEADS identifies common attributes of subdataset as specified by SQL queries and exploits this meta-information for customized, efficient and effective predictive modeling. For example, as illustrated in Figure 2(b), the SQL query specifies instances sharing attributes of gender (male), age (24), and location (NYC). LEADS leverages these propositional formulas, i.e., $\varphi$ in SQL queries, to customize the predictive model, thereby enhancing prediction accuracy for the specified subdataset.
We first propose a SQL query encoder to extract the information encapsulated in $\varphi$ into a representation vector for subsequent predictive modeling. Next, we elaborate on the model architecture and data processing flow of LEADS for SQL-aware structured data modeling. Finally, we introduce the optimization schemes, with two novel regularization terms proposed to balance the effectiveness and efficiency of the modeling process.
3.1. SQL Query Encoder
In SQL-aware predictive modeling, the WHERE clause filters tuples based on a propositional formula $\varphi$. This formula consists of predicates, each imposing a condition on a specific attribute, like ”gender = ‘M’”. These predicates are combined using logical operators such as ”AND” or ”OR” to form a propositional formula. Given the exponential number of possible predicate combinations, devising a general approach to effectively transform SQL queries into feature vectors for subsequent predictive modeling is challenging. To achieve this, we focus on individual queries, referred to as primitive SQL query. Considering a table T with $M$ attributes, the $j$-th attribute denoted as $A_{j}$, each attribute is linked to either a numerical or categorical feature. Particularly, each numerical feature needs to be converted into a corresponding categorical feature through discretization, which will be detailed subsequently. In a primitive SQL query, each attribute $A_{j}$ may be associated with zero or one predicate, with predicates across attributes conjoined using the logical operator $\land$ (AND). Technically, a predicate for the attribute $A_{j}$ in a primitive SQL query can be expressed as $P_{j}:A_{j}=a_{j}$, where $a_{j}\in\mathcal{D}_{j}\cup\{\Delta_{j}\}$, $\mathcal{D}_{j}$ represents the domain of possible values for attribute $A_{j}$, and $\Delta_{j}$ denotes a default value assigned to $A_{j}$ when it is not specified in the query. Figure 2(a) illustrates a valid primitive SQL query example, contrasting with two non-examples. Thus, the propositional formula $\varphi$ can be represented as:
$\varphi=P_{1}\land P_{2}\land\cdots\land P_{M}.$
The objective of the SQL query encoder is to generate a categorical feature vector $\mathbf{q}$ for each primitive SQL query to represent the meta-information from $\varphi$, which is achieved by concatenating the attribute values of the predicates. Formally, the feature vector of the SQL query encoding can be obtained by:
$\mathbf{q}=[q_{1},q_{2},\cdots,q_{M}]$
where $q_{j}$ is the categorical attribute value for predicate $P_{j}$. Figure 2(b) demonstrates the transformation of a primitive SQL query into a feature vector. Notably, the numerical attribute ”age” here is discretized before encoding, and any attribute $A_{i}$ without predicates is assigned the default value $\Delta_{i}$.
Discretization. Discretization is essential for encoding numerical attributes like weight or salary, as their infinite possible values make direct encoding infeasible. The discretization algorithm partitions the domain $\mathcal{D}$ of each numerical attribute into a predetermined number of bins, and then, assigns values to their respective nearest bins. This process aims to retain the key information of numerical data in the embedding space for preserving predictive capacity.
To this end, we employ a supervised discretization approach that takes into account the correlation between numerical attributes and the predictive target. This method aims to maximize information value (IV), which measures the uncertainty reduction within each bin relative to the prediction target. Higher IV values indicate a significant decrease in uncertainty, thereby preserving the predictive capacity. Specifically, we introduce OptBinning (Navas-Palencia, 2020) implementation for discretization to optimize IV while supporting constraints like the maximum bin count per attribute. The binning rule is learned on the training set and applied to the test set.
@@ 3.1. SQL Query Encoder
(b) Process of encoding a SQL query.
@@
3.2. SQL-Aware Dynamic Model Slicing
The SQL feature vector q, derived from the SQL query encoder, captures key information that can be exploited to tailor the predictive model for target subdatasets. Our objective is to construct a general model with sufficient modeling capacity and then customize this model based on q to enhance efficiency and predictive performance in SQL-aware predictive modeling. While some approaches (Von Oswald et al., 2019; Wang et al., 2010; Liu et al., 2017) modify the model’s architecture or parameters to maintain its size, they can lead to unstable predictive performance. Conversely, ensemble approaches (Ganaie et al., 2022) improve performance by stacking models, but at the cost of additional computation. To balance effectiveness and efficiency, we introduce LEADS, which scales up the modeling capacity via MoE by replicating the base models as expert models, and integrates a SQL-aware gating network based on q of SQL queries to selectively activate specific experts. The overview of LEADS is illustrated in Figure 3, and the key components following the data flow are introduced in this subsection.
Preprocessing Module There are two sets of input constructed for the SQL-aware prediction modeling given an input tuple. The first set of input is constructed for the gating network and can be uniformly represented as a categorical feature vector $\mathbf{q}$. $\mathbf{q}=[q_{1},q_{2},\ldots,q_{M}]$ comprises $M$ feature values from respective attribute fields, where numerical attributes need to be converted into categorical attributes via discretization, as described in the previous subsection. The second set is the attribute values of the input tuple $\mathbf{x}=[x_{1},x_{2},\ldots,x_{M}]$, and each attribute value $x_{i}$ can be either categorical or numerical.
For both $\mathbf{q}$ and $\mathbf{x}$, each field of attribute value $v_{i}$ ($q_{i}$/$x_{i}$) needs to be transformed into a corresponding embedding vector $\mathbf{e}_{i}$ to participate in the subsequent predictive modeling. Specifically, each categorical attribute is transformed via embedding lookup, i.e., $\mathbf{e}_{i}=\mathbf{E}_{i}[q_{i}],\mathbf{e}_{i}\in\mathbb{R}^{n_{e}}$, where $n_{e}$ is the dimension of embedding, and $\mathbf{E}_{i}$ is the embedding matrix of this categorical attribute. Note that different embedding vectors in $\mathbf{E}_{i}$ correspond to their respective attribute values. As for each numerical attribute $x_{j}$ in $\mathbf{x}$, the corresponding embedding vector is obtained by linearly scaling up a learnable embedding vector $\mathbf{\hat{e}}_{j}$ , i.e., $\mathbf{e}_{j}=x_{j}\cdot\mathbf{\hat{e}}_{j}$. In this way, we can obtain fixed-size inputs, namely embedding vectors $\mathbf{\hat{q}}=[\mathbf{q}_{1},\mathbf{q}_{2},\ldots,\mathbf{q}_{M}]$ and $\mathbf{\hat{x}}=[\mathbf{x}_{1},\mathbf{x}_{2},\ldots,\mathbf{x}_{M}]$.
@@ 3.2. SQL-Aware Dynamic Model Slicing
General Model and SQL-aware Gating Network. LEADS replicates the base model $K$ times to construct the general model, denoted as $\mathbf{\mathcal{F}}=[\mathcal{F}_{1},\mathcal{F}_{2},\ldots,\mathcal{F}_{K}]$ , where each base model is referred to as an “expert model ” in MoE. These expert models share the same architecture but learn distinct model parameters during training, which take the same input $\mathbf{\hat{x}}$ and produce different outputs that are later aggregated to form the final prediction. The output of the $i$-th expert for a given input $\mathbf{x}$ is denoted as $\mathcal{F}_{i}(\mathbf{\hat{x}})$.
@@
The SQL-aware gating network $\mathcal{G}$ takes the SQL query embedding vectors $\mathbf{\hat{q}}$ as input to produce a $K$-dimensional vector, termed the gating weight $\mathbf{w}$, with $\mathbf{w}\in\mathbb{R}^{K}$. Specifically, a two-layer multilayer perceptron (MLP) is employed as the gating network following the practice (Riquelme et al., 2021; Fedus et al., 2022; OpenAI, 2023). We concatenate all embeddings in $\mathbf{\hat{q}}$ as the input of the gating network $\widetilde{\textbf{q}}=\mathbf{q}_{1}\oplus\mathbf{q}_{2}\ldots\oplus\mathbf{q}_{M}$, where $\widetilde{\textbf{q}}\in\mathbb{R}^{M{\cdot}{n_{e}}}$, then feed $\widetilde{\mathbf{q}}$ to $\mathcal{G}$, and obtain the gating weight $\mathbf{w}$ by:
@@
$\mathbf{w}$ $=\mathcal{G}(\mathbf{s})=\mathbf{W_{2}}\mathbf{z}+\mathbf{b_{2}}$
@@
Given the gating weight $\mathbf{w}$, the $\alpha$-entmax function (Peters et al., 2019; Correia et al., 2019) is further applied to recalibrate $\mathbf{w}$ to a probability distribution. As introduced in Section 2, the hyper-parameter $\alpha$ in $\alpha$-entmax controls the level of sparsity, and a larger value of $\alpha$ sets more gating weights to zero and thus deactivates more experts for higher efficiency. The output of $\alpha$-entmax $\widetilde{\mathbf{w}}$ is:
@@
which is utilized as weights to aggregate expert outputs. The final output of the general model is a weighted average of expert outputs:
@@
where $\mathbf{\hat{y}}$ is the prediction given the input $\mathbf{x}$ and the corresponding query $\mathbf{q}$ in the SQL-aware predictive modeling.
Dynamic Model Slicing via Gating Network. Given a SQL query, all retrieved data tuples share the same recalibrated gating weight $\widetilde{\mathbf{w}}$. Notably, $\widetilde{w}_{i}=0$ in Equation 4 indicates that the $i$-th expert is not required for the current current predictive task, and thus, only a selected subset of experts $\mathcal{F}_{i}$ are activated, ensuring a more efficient inference process.
Denoting the set of indices of activated experts as $\{I_{1},I_{2},\cdots,I_{n_{o}}\}$, where $n_{o}$ is the current number of activated experts and $\widetilde{w}_{I_{j}}\neq 0,\forall j\in\{1,2,\dots,n_{o}\}$, and given the corresponding SQL feature vector $\mathbf{q}$, we index the activated experts to form a sliced model, i.e., $\mathbf{\mathcal{F}}_{\mathbf{q}}=[\mathcal{F}_{I_{1}},\mathcal{F}_{I_{2}},\cdots,\mathcal{F}_{I_{n_{o}}}]$. Thus, the final output is as follows:
@@
where the number of activated experts $n_{o}$ directly affects the effectiveness and efficiency of the sliced model. A large $n_{o}$ indicates larger model capacity while incurring higher computational overhead, and vice versa. In LEADS, $n_{o}$ is determined by the gating network based on the SQL feature vector q and the hyper-parameter $\alpha$ of the sparsemax function. Notably, instead of predefining a fixed value, $\alpha$ in $\alpha$-entmax is learnable and optimized based on the input tuples and corresponding queries during training. Subsequently, during inference, LEADS can dynamically adapt $n_{o}$ based on the current SQL query, trading off between the effectiveness and efficiency of the predictive modeling.
@@ 3.3. Optimization
Our LEADS technique can be applied to different predictive tasks by configuring a proper objective function, such as mean squared error (MSE) for regression or cross-entropy for classification. For example, in binary classification, the objective function employed is binary cross-entropy:
(6) ${\rm LogLoss}({\hat{y},y})=-\frac{1}{N}\sum_{i}^{N}\{y_{i}{\rm log}\sigma(\hat{y_{i}})+(1-y_{i}){\rm log}(1-\sigma(\hat{y_{i}}))\}$
where $\hat{y}$ is the prediction label, $y$ is the ground truth label, $N$ is the number of tuples, and $\sigma(\cdot)$ is the sigmoid function.
To make the optimization more robust and effective, we introduce two regularization terms into the main loss function. The first term is the balance loss, $\mathcal{L}_{baln}$, which is to address imbalanced expert utilization, a common issue in training MoE-based models. This imbalance arises when the gating network $\mathcal{G}$ disproportionately favor a few experts, which skews the training dynamics. As a consequence, preferred experts become overutilized while others remain underutilized, which compromises the MoE capacity and degrades overall model performance.
Let $\mathbf{X}$ denote a mini-batch of training instances with $n_{b}$ tuples, and $\widetilde{\mathbf{W}}=[\widetilde{\mathbf{w}}_{1},\widetilde{\mathbf{w}}_{2},\cdots,\widetilde{\mathbf{w}}_{n_{b}}]$ represent the recalibrated gating weights of $\mathbf{X}$. Here, $\widetilde{w}_{ij}$ is the $j$-th weight of $\widetilde{\mathbf{w}}_{i}$. $\mathcal{L}_{baln}$ is defined as follows:
@@
where $\mathbb{E}(\mathbf{\Phi})=\frac{1}{K}\sum_{j=1}^{K}\phi_{j}$. This balance loss term promotes a uniform distribution of weights across experts in the mini-batch, encouraging equal contribution from each expert. However, this term empirically tends to activate a large number of experts due to its drive for balance, which is contrary to the intent of sparse softmax to maintain sparsity. To counterbalance this and enhance computational efficiency, we further introduce an additional term, the sparsity loss term, $\mathcal{L}_{sprs}$:
@@
which encourages the gating network to allocate higher weights on a select few experts while assigning minimal or zero weights to others. Both loss terms are weighted by their respective regularization coefficient, $\lambda_{1}$ and $\lambda_{2}$, and incorporated into the overall objective function:
(9) ${\rm Loss}(\hat{\mathbf{y}},{\mathbf{y}})={\rm LogLoss}(\hat{\mathbf{y}},{\mathbf{y}})+\lambda_{1}L_{baln}+\lambda_{2}L_{sprs}.$
With this objective function, LEADS can then be optimized effectively using gradient-based optimizers, e.g., SGD  (Amari, 1993) or Adam (Kingma and Ba, 2014).
Figure 4. Breakdown of inference response time for the Inference-Decouple Strategy.